# Invariant Coordinate Selection and Fisher discriminant subspace beyond the case of two groups

### v1

@@ 6 Empirical study
In Section 3, we derive a theoretical understanding of the behavior of the eigenvalues of $\mathrm{V}_{1}^{-1}\mathrm{V}_{2}$ in the context of a mixture of Dirac distributions (model (5)). Firstly, for a mixture model with full rank group centers matrix and in the absence of within-group variability, the eigenvalues of $\mathrm{V}_{1}^{-1}\mathrm{V}_{2}$, depend only on the group proportions (see Proposition 1). Secondly, in Subsection 4.2, the thresholds for the proportion of groups that result in an eigenvalue transitioning from less than the eigenvalue of multiplicity $p-q$ to greater than it, have been identified for different number of groups. Both analyses are restricted to the case where the dimension of the FDS is $q=k-1$. To confirm those results in a more general context but still with $q=k-1$, we perform simulations of a mixture of Gaussian distributions, including within-group variability and noise, and we also focus on different scatter pairs. Subsection 6.1 describes the simulations settings, Subsection 6.2 studies the eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ when group centers vary in the presence of within-group variability, while Subsection 6.3 focuses on the behavior of eigenvalues for different scatter pairs.
@@ 6.1 Simulation design
$X\sim\sum_{j=1}^{k}\alpha_{j}\,{\cal N}(t_{j},I_{p}),$ (8)
where $\alpha_{j}>0$ are mixture weights such that $\sum_{j=1}^{k}\alpha_{j}=1$, for $j\in\{1,\ldots,k\}$, $t_{1}=0$, and $t_{l+1}=\delta e_{l}$, for $l\in\{1,\ldots,k-1\}$, $e_{l}$ is a $p$-dimensional vector with one in the $l$-th coordinate and zero elsewhere, for different numbers of clusters $k\in\{2,3,5,10\}$, on $p=5k$ variables. In this setting, the cluster structure lies in a low-dimensional subspace of dimension $q=k-1$. We consider 17 scenarios of group proportions ($\alpha_{j}$, $j\in\{1,\ldots,k\}$):
@@
We consider $\delta\in\{1,5,10,50,100\}$ for the study on different group centers in Subsection 6.2. For the study on different scatters in Subsection 6.3, we fix $\delta=10$ and we focus on the following scatter pairs: $\mathrm{COV}-\mathrm{COV}_{4}$, $\mathrm{COVAXIS}-\mathrm{COV}$, $\mathrm{TCOV}-\mathrm{COV}$ and $\mathrm{MCD}_{\tau}-\mathrm{COV}$ with $\tau\in\{0.25,0.5,0.75\}$. $\mathrm{COVAXIS}$ is a one-step M-estimator using the inverse weight function of the squared Mahalanobis distance, $\mathrm{TCOV}$ is also a one-step M-estimator but with weights based on pairwise Mahalanobis distances and $\mathrm{MCD}_{\tau}$ are the (raw) minimum covariance determinant estimators based on $n_{\tau}=\lceil\tau n\rceil$ observations for which the sample covariance matrix has the smallest determinant. See Subsection 3.1 of Alfons et al. (2024) for details on those scatter matrices. Note that we follow Alfons et al. (2024) and take $\mathrm{V}_{1}$ being more robust than $\mathrm{V}_{2}$. In addition, $\mathrm{MCD}_{\tau}$ are computed using the FAST-MCD algorithm of (Rousseeuw and Driessen, 1999). For the selection of components, we choose the med criterion as introduced in Alfons et al. (2024), which selects the $k-1$ components whose eigenvalues deviate the most from the median of all eigenvalues. This test relies on the assumption that the dimension of interest $q$ is low compared to the number of variables $p$: $q\leq p/2$, which implies that the majority of the eigenvalues should be equal, which is met in our context. We simulate 50 data sets for each of the different settings.
@@
As proven in Proposition 1, theoretically, for a mixture model with full rank group centers matrix and in the absence of within-group variability, the eigenvalues of $\mathrm{V}_{1}^{-1}\mathrm{V}_{2}$ depend only on the group proportions $\alpha_{j}$, $j\in\{1,\ldots,k\}$ of the mixture components. Here, instead of computing the eigenvalues theoretically, we simulate a mixture of Gaussian distributions with some within-group variability to evaluate the sensitivity of this proposition and we compare with the results in Figure 7 for $\mathrm{COV}^{-1}\mathrm{COV}_{4}$. In Figure 8, we display the boxplots of the $k-1$ first and the $k-1$ last eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ over 50 replications when the group centers vary in the presence of within-group variability. Grey-shaded areas expose the eigenvalues which are theoretically different from one. The cases for which the eigenvalue one has multiplicity greater than $p-q$ are identified by the red-shaded areas.
@@
Contrary to Figure 7, we can see some variability for the eigenvalues over the replications and between the different group centers, determined by the values of $\delta$. In addition, given the presence of within-group variability and noise with $p>q=k-1$, none of the eigenvalues are strictly equal to one. However, the eigenvalues of interest for highlighting the groups’ structure (in grey-shaded areas), are still easily identifiable as the ones further away from one. For example, for the first row with scenarios of a mixture of two components, only the first or the last eigenvalue should be different from one. More specifically, as illustrated in Subsection 4.2, for a mixture with group proportion of $50-50$, only the last eigenvalue is different from one, as we can see in the first subplot. If one group proportion decreases until being below 21% then it is now the first eigenvalue which is different from one. For this threshold of 21%, all eigenvalues are equal to one as highlighted by the red-shaded area. We can identify the same behavior when the number of groups increases. For example, in the second row, if all the three group proportions are “large enough” then the last two eigenvalues are different from one. When only one group proportion is “low”, as in the scenario $10-40-50$, the first and the last components are different from one. When two group proportions are “low”, like in the scenario $10-10-80$, then only the first two components are different from one. This pattern repeats itself for mixtures of 5 or 10 groups as illustrated in the third and fourth rows. The thresholds are clearly identifiable when one group proportion is equal to: 18% for 3 groups, 14% for 5 groups and 8% for 10 groups and confirms the ones identified in Subsection 4.2 in case of absence of within-group variability.
@@
It is also important to note that the variability of the eigenvalues between group centers, i.e for different values of $\delta$, is smaller for the eigenvalues which are theoretically equal to one (in white or red-shaded areas) no matter the value of $\delta$. This supports evidence for some stability of the eigenvalues associated with no structure of the data, no matter the group centers. Clearly, at the thresholds, identified by red-shaded areas, there is almost no variability of the eigenvalues, as we can see in the first row, in the well-known case of a mixture of two components with group proportions $21-79$. For the eigenvalues of interest (in grey-shaded areas), some variability is visible but they seem to quickly converge as $\delta$ increases. In addition, the eigenvalues are clearly different from one as soon as $\delta\geq 5$. For $\delta=1$, the ratio signal over noise is too low and so, all eigenvalues are close to one. On the contrary, if $\delta\geq 50$, the ratio signal over noise is high and the noise is almost null so, we are almost in the same context as for the Proposition 1, and the eigenvalues with $\delta=50$ or $\delta=100$ are almost equal. In fact, it appears that the eigenvalues converge more slowly in the presence of a low group proportion, as observed in scenarios where at least one group has a proportion of 5% or 10%. In this context, the eigenvalues are also higher in absolute value for any $\delta$ so it is one of the favorable cases for ICS since it is easy to identify it is different from one. So, for the next study, we consider that the group centers have a negligible effect, and we focus on the case $\delta=10$.
@@
6.3 $\mathrm{V}_{1}^{-1}\mathrm{V}_{2}$ eigenvalues for different scatter pairs in the presence of within-group variability and for $q=k-1$
@@
In this section, we extend our study to the behavior of the eigenvalues of $\mathrm{V}_{1}^{-1}\mathrm{V}_{2}$ for different affine equivariant scatter pairs and not only to $\mathrm{COV}-\mathrm{COV}_{4}$. In such a context, and in the absence of within-group variability, only the group proportions should impact the eigenvalues. Here, we introduce some within-group variability and we analyze if the general behavior for the eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ is the same for different scatter pairs. More specifically, we want to know if the thresholds are linked to the same group proportions and if only “low” group proportions are associated with first eigenvalues being different from each other. Deriving theoretical eigenvalues for more complex scatter pairs is a difficult task so, we focus on the analysis of the invariant components instead. Indeed, to highlight the data structure, we have to project the data onto the subspace spanned by the eigenvectors associated with eigenvalues which are not equal to each other. So we can hypothesize that the invariant components to select are the ones associated with the eigenvalues which are not equal to each other. To identify the components of interest, we followed the recommendations of Alfons et al. (2024) and we chose the $k-1$ invariant components based on the med criterion, i.e the ones associated with eigenvalues which deviate the most from the median of all eigenvalues.
@@
In Figure 9, we display heatmaps of the percentage of selections for the $k-1$ first and the $k-1$ last invariant components of $\mathrm{V}^{-1}\mathrm{V}_{2}$ by the med criterion, for different numbers of groups, for group proportions and for different scatter pairs. If the value is of 100%, it indicates that the value of the eigenvalue associated with this component is stable and different from each others. If the percentage is less than 100%, it means that over the replications, the criterion has not chosen consistently the same eigenvalue which suggests that the eigenvalue might be not significantly different from the other ones. For example, in the first subplot, for the scatter pair $\mathrm{COV}-\mathrm{COV}_{4}$, the value is 100% for the case of a mixture of two Gaussian distributions with group proportions of 10-90. We can deduce that the first component IC1 is always chosen over the 50 replications. This is in line with the previous conclusion which states that if one group proportion is “low" then the first eigenvalue is different from one. In case of 50-50, it is always the last one which is chosen and of interest. For group proportions in between, the choice is not as clear mainly for the scenario 21-79. In this case, IC1 is selected instead of IC10 only in 75% of cases. When the selection is not clear, this is a sign that the eigenvalues are not as far away from the others and it might indicate a threshold. Precisely, this result confirms the theory recalled in Subsection 2.3 that, when a group proportion is equal to 21% and in presence of two groups, then all the eigenvalues are equal to one with $\mathrm{COV}-\mathrm{COV}_{4}$. In practice, the group proportion associated with the threshold is not as precise as in the theoretical situation since, even when a group proportion is around 30%, it can be difficult to have a clear separation between eigenvalues in 15% of cases.
@@
Fig. 9: Heatmaps of the percentage of selection for the $k-1$ first and $k-1$ last selected invariant components of $\mathrm{V}^{-1}\mathrm{V}_{2}$ by the med criterion, in the presence of within-group variability, for $q=k-1$, $p=k/0.2$, $n=1000$, different values of $k$ over 50 replications. Results for different numbers of groups are shown in separate rows, for different group proportions in the y-axis and for different scatter pairs in different panels.
@@
Overall, for $\mathrm{COV}-\mathrm{COV}_{4}$ we can corroborate the same thresholds as mentioned in the previous section: approximately 21% for 2 groups, 18% for 3 groups, 14% for 5 groups and 8% for 10 groups. However, because we infer the behavior of the eigenvalues based on the selection of components and we are in the context of within-group variability, noise and simulated data, the values of the thresholds are not as precise as in the theoretical cases illustrated in Figure 7. Nevertheless, this analysis globally confirms the results already observed and demonstrated in Subsection 4.2. It is also quite easy to visually identify on which components the structure of interest relies depending on the group proportions and to extend our understanding of such behavior for the other scatter pairs.

### v2

@@ 6 Empirical study
In Section 3, we derive a theoretical understanding of the behavior of the eigenvalues of $\mathbf{V}_{1}^{-1}\mathbf{V}_{2}$ in the context of a mixture of Dirac distributions (model (5)). Firstly, for a mixture model with full rank group centers matrix and in the absence of within-group variability, the eigenvalues of $\mathbf{V}_{1}^{-1}\mathbf{V}_{2}$, depend only on the group proportions (see Proposition 1). Secondly, in Subsection 4.2, the thresholds for the proportion of groups that result in an eigenvalue transitioning from less than the eigenvalue of multiplicity $p-q$ to greater than it, have been identified for different number of groups. Both analyses are restricted to the case where the dimension of the FDS is $q=k-1$. To confirm those results in a more general context but still with $q=k-1$, we perform simulations of a mixture of Gaussian distributions, including within-group variability and noise, and we also focus on different scatter pairs. Subsection 6.1 describes the simulations settings, Subsection 6.2 studies the eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ when group centers vary in the presence of within-group variability, while Subsection 6.3 focuses on the behavior of eigenvalues for different scatter pairs.
@@ 6.1 Simulation design
$\mathbf{X}\sim\sum_{j=1}^{k}\alpha_{j}\,{\cal N}(\mathbf{t}_{j},\mathbf{I}_{p}),$
where $\alpha_{j}>0$ are mixture weights such that $\sum_{j=1}^{k}\alpha_{j}=1$, for $j\in\{1,\ldots,k\}$, $\mathbf{t}_{1}=0$, and $\mathbf{t}_{l+1}=\delta\mathbf{e}_{l}$, for $l\in\{1,\ldots,k-1\}$, $\mathbf{e}_{l}$ is a $p$-dimensional vector with one in the $l$-th coordinate and zero elsewhere, for different number of clusters $k\in\{2,3,5,10\}$, on $p=5k$ variables. In this setting, the cluster structure lies in a low-dimensional subspace of dimension $q=k-1$. We consider 17 scenarios of group proportions ($\alpha_{j}$, $j\in\{1,\ldots,k\}$):
@@
We consider $\delta\in\{1,5,10,50,100\}$ for the study on different group centers in Subsection 6.2. For the study on different scatters in Subsection 6.3, we fix $\delta=10$ and we focus on the following scatter pairs: $\mathrm{COV}-\mathrm{COV}_{4}$, $\mathrm{COVAXIS}-\mathrm{COV}$, $\mathrm{TCOV}-\mathrm{COV}$ and $\mathrm{MCD}_{\tau}-\mathrm{COV}$ with $\tau\in\{0.25,0.5,0.75\}$. $\mathrm{COVAXIS}$ is a one-step M-estimator using the inverse weight function of the squared Mahalanobis distance, $\mathrm{TCOV}$ is also a one-step M-estimator but with weights based on pairwise Mahalanobis distances and $\mathrm{MCD}_{\tau}$ are the (raw) minimum covariance determinant estimators based on $n_{\tau}=\lceil\tau n\rceil$ observations for which the sample covariance matrix has the smallest determinant. See Subsection 3.1 of Alfons et al. (2024) for details on those scatter matrices. Note that we follow Alfons et al. (2024) and take $\mathbf{V}_{1}$ being more robust than $\mathbf{V}_{2}$. In addition, $\mathrm{MCD}_{\tau}$ are computed using the FAST-MCD algorithm of (Rousseeuw and Driessen, 1999). For the selection of components, we choose the med criterion as introduced in Alfons et al. (2024), which selects the $k-1$ components whose eigenvalues deviate the most from the median of all eigenvalues. This test relies on the assumption that the dimension of interest $q$ is low compared to the number of variables $p$: $q\leq p/2$, which implies that the majority of the eigenvalues should be equal, which is met in our context. We simulate 50 data sets for each of the different settings.
@@
As proven in Proposition 1, theoretically, for a mixture model with full rank group centers matrix and in the absence of within-group variability, the eigenvalues of $\mathbf{V}_{1}^{-1}\mathbf{V}_{2}$ depend only on the group proportions $\alpha_{j}$, $j\in\{1,\ldots,k\}$ of the mixture components. Here, instead of computing the eigenvalues theoretically, we simulate a mixture of Gaussian distributions with some within-group variability to evaluate the sensitivity of this proposition and we compare with the results in Fig. 7 for $\mathrm{COV}^{-1}\mathrm{COV}_{4}$. In Fig. 8, we display the boxplots of the $k-1$ first and the $k-1$ last eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ over 50 replications when the group centers vary in the presence of within-group variability. Grey-shaded areas expose the eigenvalues which are theoretically different from one. The cases for which the eigenvalue one has multiplicity greater than $p-q$ are identified by the red-shaded areas.
@@
Contrary to Fig. 7, we can see some variability for the eigenvalues over the replications and between the different group centers, determined by the values of $\delta$. In addition, given the presence of within-group variability and noise with $p>q=k-1$, none of the eigenvalues are strictly equal to one. However, the eigenvalues of interest for highlighting the groups’ structure (in grey-shaded areas) are still easily identifiable as the ones further away from one. For example, for the first row with scenarios of a mixture of two components, only the first or the last eigenvalue should be different from one. More specifically, as illustrated in Subsection 4.2, for a mixture with group proportion of $50-50$, only the last eigenvalue is different from one, as we can see in the first subplot. If one group proportion decreases until being below 21% then it is now the first eigenvalue which is different from one. For this threshold of 21%, all eigenvalues are equal to one as highlighted by the red-shaded area. We can identify the same behavior when the number of groups increases. For example, in the second row, if all the three group proportions are “large enough” then the last two eigenvalues are different from one. When only one group proportion is “low”, as in the scenario $10-40-50$, the first and the last components are different from one. When two group proportions are “low”, like in the scenario $10-10-80$, then only the first two components are different from one. This pattern repeats itself for mixtures of 5 or 10 groups as illustrated in the third and fourth rows. The thresholds are clearly identifiable and equal to 18% for 3 groups, 14% for 5 groups and 8% for 10 groups, confirming those identified in Subsection 4.2 in the absence of within-group variability.
@@
It is also important to note that the variability of the eigenvalues between group centers, i.e., for different values of $\delta$, is smaller for the eigenvalues which are theoretically equal to one (in white or red-shaded areas) no matter the value of $\delta$. This supports evidence for some stability of the eigenvalues associated with no structure of the data, no matter the group centers. Clearly, at the thresholds, identified by red-shaded areas, there is almost no variability of the eigenvalues, as we can see in the first row, in the well-known case of a mixture of two components with group proportions $21-79$. For the eigenvalues of interest (in grey-shaded areas), some variability is visible but they seem to quickly converge as $\delta$ increases. In addition, the eigenvalues are clearly different from one as soon as $\delta\geq 5$. For $\delta=1$, the ratio signal over noise is too low and so, all eigenvalues are close to one. On the contrary, if $\delta\geq 50$, the ratio signal over noise is high and the noise is almost null so, we are almost in the same context as for the Proposition 1, and the eigenvalues with $\delta=50$ or $\delta=100$ are almost equal. In fact, it appears that the eigenvalues converge more slowly in the presence of a low group proportion, as observed in scenarios where at least one group has a proportion of 5% or 10%. In this context, the eigenvalues are also higher in absolute value for any $\delta$ so it is one of the favorable cases for ICS since it is easy to identify it is different from one. So, for the next study, we consider that the group centers have a negligible effect, and we focus on the case $\delta=10$.
@@
6.3 $\mathbf{V}_{1}^{-1}\mathbf{V}_{2}$ eigenvalues for different scatter pairs in the presence of within-group variability and for $q=k-1$
@@
In this section, we extend our study to the behavior of the eigenvalues of $\mathbf{V}_{1}^{-1}\mathbf{V}_{2}$ for different affine equivariant scatter pairs and not only to $\mathrm{COV}-\mathrm{COV}_{4}$. In such a context, and in the absence of within-group variability, only the group proportions should impact the eigenvalues. Here, we introduce some within-group variability and we analyze if the general behavior for the eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ is the same for different scatter pairs. More specifically, we want to know if the thresholds are linked to the same group proportions and if only “low” group proportions are associated with first eigenvalues being different from each other. Deriving theoretical eigenvalues for more complex scatter pairs is a difficult task so, we focus on the analysis of the invariant components instead. Indeed, to highlight the data structure, we have to project the data onto the subspace spanned by the eigenvectors associated with eigenvalues which are not equal to each other. So we can hypothesize that the invariant components to select are the ones associated with the eigenvalues which are not equal to each other. To identify the components of interest, we followed the recommendations of Alfons et al. (2024) and we chose the $k-1$ invariant components based on the med criterion, i.e., the ones associated with eigenvalues which deviate the most from the median of all eigenvalues.
@@
In Fig. 9, we display heatmaps of the percentage of selections for the $k-1$ first and the $k-1$ last invariant components of $\mathbf{V}^{-1}\mathbf{V}_{2}$ by the med criterion, for different number of groups, for group proportions and for different scatter pairs. If the value is 100%, it indicates that the eigenvalue associated with this component is stable and different from the others. If the percentage is less than 100%, it means that, over the replications, the criterion has not consistently chosen the same eigenvalue, suggesting that the eigenvalue might not be significantly different from the others. For example, in the first subplot, for the scatter pair $\mathrm{COV}-\mathrm{COV}_{4}$, the value is 100% for the case of a mixture of two Gaussian distributions with group proportions of 10-90. We can deduce that the first component IC1 is always chosen over the 50 replications. This is in line with the previous conclusion which states that if one group proportion is “low" then the first eigenvalue is different from one. In case of 50-50, it is always the last one which is chosen and of interest. For group proportions in between, the choice is not as clear mainly for the scenario 21-79. In this case, IC1 is selected instead of IC10 only in 75% of cases. When the selection is not clear, this is a sign that the eigenvalues are not as far away from the others and it might indicate a threshold. Precisely, this result confirms the theory recalled in Subsection 2.3 that, when a group proportion is equal to 21% and in presence of two groups, then all the eigenvalues are equal to one with $\mathrm{COV}-\mathrm{COV}_{4}$. In practice, the group proportion associated with the threshold is not as precise as in the theoretical situation since, even when a group proportion is around 30%, it can be difficult to have a clear separation between eigenvalues in 15% of cases.
@@
Fig. 9: Heatmaps of the percentage of selection for the $k-1$ first and $k-1$ last selected invariant components of $\mathbf{V}^{-1}\mathbf{V}_{2}$ by the med criterion, in the presence of within-group variability, for $q=k-1$, $p=k/0.2$, $n=1000$, different values of $k$ over 50 replications. Results for different number of groups are shown in separate rows, for different group proportions in the y-axis and for different scatter pairs in different panels.
@@
Overall, for $\mathrm{COV}-\mathrm{COV}_{4}$ we can corroborate the same thresholds as mentioned in the previous section: approximately 21% for 2 groups, 18% for 3 groups, 14% for 5 groups and 8% for 10 groups. However, because we infer the behavior of the eigenvalues based on the selection of components and we are in the context of within-group variability, noise and simulated data, the values of the thresholds are not as precise as in the theoretical cases illustrated in Fig. 7. Nevertheless, this analysis globally confirms the results already observed and demonstrated in Subsection 4.2. It is also quite easy to visually identify on which components the structure of interest relies depending on the group proportions and to extend our understanding of such behavior for the other scatter pairs.
@@
7 Empirical application
To extend our findings beyond theoretical distributions, we analyze the Olympic decathlon dataset from the 2016 Rio de Janeiro Games (Franceschini and Loperfido, 2017), which contains scores across 10 events for 23 athletes. While PCA only identifies two outlying athletes (Saluri and Nakamura) in its first two components, Loperfido (2018) showed that skewness-based projection pursuit reveals additional structure. Notably, it detects the athleteTaiwo, who despite ranking 11th overall, displayed unusual performance patterns – achieving both the highest High Jump score and nearly the lowest Javelin score.
As shown in Fig. 10 (first row), ICS with $\mathrm{COV}-\mathrm{COV}_{4}$ identifies Saluri, Taiwo, and Nakamura on IC1, IC2, and IC3 respectively, complementing both PP and PCA findings by detecting all three outliers. This matches our theoretical results from Sections 3, 4, and 6, where we showed that eigenvalue transitions depend solely on group proportions. For $\mathrm{COV}-\mathrm{COV}_{4}$, the threshold decreases from 21% (two groups) to 8% (ten groups) in mixture distributions. Viewing the data as a mixture model with three outliers (Taiwo, Nakamura, and Saluri, each  4.3%) and a majority group (86.9%) explains why these athletes are detected on the first invariant components, consistent with common mixture modeling of outliers (Archimbaud et al. (2018), Hou and Wentzell (2014)).
To further analyze eigenvalue behavior, let us add a new athlete to the dataset who behaves similarly to Saluri, the lowest-ranked athlete. We apply a uniform $U_{[-20,20]}$ noise to each of Saluri’s scores. This yields mixture weights of approximately $\alpha_{Saluri}\approx 8.3\%$, $\alpha_{Taiwo}=\alpha_{Nakumara}\approx 4.1\%$, and $\alpha_{majority}\approx 83.3\%$. As shown in the second row of Fig. 10, Saluri is no longer an outlier, while Taiwo appears on the first component and Nakamura on the second. Next, we expand the Saluri-based cluster to weights of $\alpha_{Saluri}\approx 12\%$, $\alpha_{Taiwo}=\alpha_{Nakumara}\approx 4\%$, and $\alpha_{majority}\approx 80\%$ by adding another synthetic observation. In the last row of Fig. 10, Taiwo and Nakamura still appear as outliers on the first two components, but the Saluri-based cluster now stands out on the last component. This supports the theory that at a given threshold, one eigenvalue transitions from being smaller to larger than the eigenvalue of multiplicity $p-q$. For this dataset, the transition happens when a group reaches roughly 8%: below this, it appears on the first component, and above this, on the last. Notably, this 8% threshold aligns with an assumption of 10 groups rather than 4 (see Section 4), suggesting that the data may naturally form 10 clusters, matching its 10 variables. Thus, even with unknown data distribution, the theoretical results hold and may provide insights into the number of clusters.
Fig. 10: Scatterplots of the second, third and last ICS components versus the first component with $\mathrm{COV}-\mathrm{COV}_{4}$ on the decathlon data set. Results for different modified data sets are shown in separate rows. The names of the athlete are preceded by their rank. The colors identify the athletes coming from the same cluster.